2.A horizontal pipe is 2 cm in radius at one end and increases in size so that it is 5 cm in radius at


the other end. Water is pumped into the small end of the pipe at 10 m/s and a pressure of


300 kPa (kilopascal). Find the


a.) speed and


b.) pressure of the water at the pipe’s large end.

1
Expert's answer
2021-12-12T16:43:09-0500

Explanations & Calculations


  • Hello Gingging,


  • This whole thing deal with the Bernoulli's law and the concept of volumetric flow rate.
  • Since the water is incompressible, the volumetric flow rate ("\small Q = Av" ) is constant throughout the pipe.

a)

  • Applying volumetric flow rate to the input and output ends,

"\qquad\qquad
\begin{aligned}
\small Q=A_1v_1&=\small A_2v_2\\
\small v_2&=\small \frac{A_1}{A_2}.v_1\\
&=\small \frac{\pi r_1^2}{\pi r_2^2}.v_1\\
&=\small \Big(\frac{r_1}{r_2}\Big)^2.v_1
\end{aligned}"

  • Here "\small r_1\,\&\,v_1" refer to the input side and the out put side follows the notation accordingly.


b)

  • For this part, you can use the Bernoulli's equation — "\small P+\frac{1}{2}\rho v^2+\rho g h=\text{constant}" — for the input and out put conditions.
  • Since the pipe is level throughout, any graduation is neglected, so that "\small h = 0" .
  • By now you have found the speed at the output ("\small v_2" ) which is needed for this step.

"\qquad\qquad
\begin{aligned}
\small P_1+\frac{1}{2}\rho v_1^2 &=\small P_2+\frac{1}{2}\rho v_2^2\\
\small P_2 &=\small P_1+\frac{1}{2}\rho(v_1^2-v_2^2)
\end{aligned}"


  • Now you can give it a try substituting the values accordingly and getting the answers.
  • use the units correctly, when the lengths are input in meters, the speeds are obtained in meters per second.
  • Let me know in comments if you find any difficulty.